A basketball of mass 0.0400 kg is dropped from a height of 5.00 m to the


ground and bounces back to a height of 3.00 m.


a. On its way down, how much potential energy does the ball lose?


b. On its way back, how much potential energy does the ball regain

1
Expert's answer
2021-11-14T17:26:09-0500

Mass of basketball = 0.0400 kg

g =<acceleration due to gravity = 9.8m/s^2

a) on its way down

Change in height = final height - initial height

= 0/00m - 5.00m = -5.00m

Loss in potential energy in way down

= Mass × acceleration due to gravity × decrease in height

"= 0.0400 kg \u00d7 9.8 m\/s^2 \u00d7-5.00m =-1.96joules"

Therefore, in its way down ball loses 1.96 joules potential energy.


b) gain in potential energy in way back

= Mass × acceleration due to gravity × increase in height

"=0.0400kg \u00d7 9.8m\/s^2 \u00d7 (+3.00m) =+1.176 joules."

On its way back ball gains potential energy = +1.176 joules